# A Computational Tsirelson’s Theorem for the Value of Compiled XOR Games

### v2

@@ 4.2 Bounding the compiled value using cryptographic pseudo-expectations
    $\tilde{\operatorname*{\mathbb{E}}}[\cdot]$ is linear,
@@
$\tilde{\mathbb{E}}_{S}\left[a_{x}a_{z}\right]$ =$=\begin{cases}1,\text{ if $x=z$}\\ 0,\text{ otherwise}\end{cases}$ (4.3)
@@ 4.3 A nice SOS certificate for any XOR game
$=\sum_{x\in I_{A},y\in I_{B}}\pi(x,y)\sum_{a,b}c_{xy}(-1)^{a+b}\operatorname*{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x)}\sum_{\alpha;\mathrm{Dec}(\alpha)=a}\langle\psi|(A_{c\alpha})^{\ast}B_{yb}A_{c\alpha}|\psi\rangle$
@@
$-\sum_{x\in I_{A}}r_{x}\left(\hat{b}_{x}^{2}-\mathds{1}\right)=\beta\mathds{1}-\sum_{x\in I_{A}}r_{x}\hat{b}_{x}^{2}=\sum_{y\in I_{B}}c_{y}b_{y}^{2}-\sum_{x\in I_{A}}r_{x}\hat{b}_{x}^{2}$ (4.13)
@@
In addition, the matrix $\Delta(\mathbf{r})$ is invertible since each $r_{x}$ is strictly positive. Hence, by the Schur complement condition for positivity, we obtain
@@ 5 Magic square game
by using $(x+y)^{2}\leq 2x^{2}+2y^{2}$, Lemmas 5.6 and 5.5. Similarly, we obtain
@@ Acknowledgments
GM, SS and MW acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2092 CASA - 390781972. GM also acknowledges support by the European Research Council through an ERC Starting Grant (Grant agreement No. 101077455, ObfusQation). MW also acknowledges support by the European Research Council through an ERC Starting Grant (Grant agreement No. 101040907, SYMOPTIC), by the NWO through grant OCENW.KLEIN.267, and by the BMBF through project Quantum Methods and Benchmarks for Resource Allocation (QuBRA). MW acknowledges the Simons Institute for the Theory of Computing at UC Berkeley for its hospitality and support while this paper was completed. AM is supported by NSERC Alliance Consortia Quantum grants, reference number number: ALLRP 578455 - 22. AN acknowledges SLMATH for hospitality and financial support during the MIP* = RE hot topics workshop. AN and TZ acknowledge the Simons Institute at UC Berkeley, on whose whiteboard a crucial calculation was performed.

### v3

@@ 3.2 Bounding the compiled value using cryptographic pseudo-expectations
    $\tilde{\operatorname*{\mathbb{E}}}$$\cdot$] is linear,
@@
$\tilde{\mathbb{E}}_{S}\left[a_{x}a_{z}\right]$ $=\begin{cases}1,\text{ if $x=z$}\\ 0,\text{ otherwise}\end{cases}$ (3.3)
@@ 3.3 A nice SOS certificate for any XOR game
$=\sum_{x\in I_{A},y\in I_{B}}\pi(x,y)\sum_{a,b}(-1)^{g(x,y)}(-1)^{a+b}\operatorname*{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x)}\sum_{\alpha;\mathrm{Dec}(\alpha)=a}\langle\psi|(A_{c\alpha})^{\ast}B_{yb}A_{c\alpha}|\psi\rangle$
@@
$\sum_{x\in I_{A}}r_{x}\left(\hat{b}_{x}^{2}-\mathds{1}\right)=\beta\mathds{1}-\sum_{x\in I_{A}}r_{x}\hat{b}_{x}^{2}=\sum_{y\in I_{B}}c_{y}b_{y}^{2}-\sum_{x\in I_{A}}r_{x}\hat{b}_{x}^{2}$ (3.13)
@@
In addition, the matrix $\Delta(\mathbf{r})$ is invertible since each $r_{x}$ is strictly positive. Hence, by the Schur complement [12, Theorem 1.12 (b)] condition for positivity, we obtain
@@
Together, Theorem 2.14 and Theorem 3.1 imply that the “asymptotic” quantum value of a compiled XOR game is equal to the quantum value of the underlying XOR game. That is, if $S$ is any quantum strategy for the compiled game, then
and conversely, there always exists a strategy such that
4 Self-testing results for compiled XOR games
see [24, Section 3]. Note that the quantity $\varepsilon$ refers to the winning probability, whereas in [24, Section 3], the notion of $\varepsilon$-optimality is defined with respect to the bias.
Definition 4.1.
Lemma 4.2.
[…]
where $\varepsilon^{\prime}=2\varepsilon+\eta^{\prime}(\lambda)$.
Proof.
Since $S$ is $\varepsilon$-optimal, and using Eqs. A.6 and A.8 in the proof of Theorem 3.1 (with $x_{0}=z$), we see that
$\leq\beta^{*}_{q}(\mathcal{G})+\eta^{\prime}(\lambda),$ (4.2)
From this we can draw two conclusions:
First, since the second inequality in Eq. 4.2 holds for each $x\in I_{A}$ (by Lemma A.2), we obtain that
where $\varepsilon^{\prime}:=2\varepsilon+\eta^{\prime}(\lambda)$.
From this, it follows that
[…]
where we used that $r_{x}\leq 1$ in the last step.
Second, from Eq. 4.2 we see that the vector strategy $\{u_{x},v_{y}\}$ is $\varepsilon^{\prime}$-optimal. Thus it follows from Eqs. 4.3 and 4.1 that
[…]
for all $x\in I_{A}$. The lemma follows from this and Eq. 4.4 for $x=z$. ∎
Theorem 4.3.
[…]
is a constant that depends only on the XOR game.
Proof.
We deduce
[…]
from Lemma 4.2. ∎
Then, Theorem 4.3 yields
5 Parallel repetition of XOR games
We show an upper bound on the quantum value of the compiled parallel repetition of XOR games. Throughout this section, we will treat the number of parallel repetitions $n$ as fixed (in other words, as a constant in the security parameter). However, we remark that a similar analysis can be used to establish the soundness of any polynomial number $n$ of parallel repetitions, at the cost of a sub-exponential security guarantee for the QHE. A similar tradeoff was also shown in [14].
Definition 5.1.
While the conjunction of XOR games is not an XOR game, we can make the following definition. We use the notation $n]=\{1,\dots,n\}$.
Definition 5.2.
We first discuss the results of [7] for nonlocal games. Clearly, any quantum strategy $S$ for $\wedge_{i=1}^{n}\mathcal{G}_{i}$ determines one for $\oplus_{i\in M}\mathcal{G}_{i}$, for any subset $M\subseteq[n]$, as follows: Alice and Bob first proceed as in strategy $S$ to obtain $a,b\in\{0,1\}^{n}$. They then output $\oplus_{i=1}^{n}a_{i}$ and $\oplus_{i=1}^{n}b_{i}$, respectively. Then one has the following relation between the quantum value of the parallel repetition and the bias of the sum games:
Lemma 5.3 ([7, Lemma 8]).
It follows from this and the nontrivial result that the bias of XOR games is multiplicative with respect to taking sums [7, Theorem 1] that one has the upper bound
The converse inequality holds for any nonlocal game, since the players can always play optimal strategies for the individual games $G_{i}$ in parallel. This establishes the parallel repetition theorem for XOR games [7, Theorem 2]:
We now use a similar argument to obtain a repetition theorem in the compiled setting. As above, we first argue that any quantum strategy $S$ for the compiled parallel repetition $(\wedge_{i=1}^{n}\mathcal{G}_{i})_{\mathrm{comp}}$ determines a quantum strategy $S_{M}$ for the compile sum games $(\oplus_{i\in M}\mathcal{G}_{i})_{\mathrm{comp}}$, for any subset $M\subseteq[n]$. To this end, we modify step 2 such that if $S$ were to output $\alpha$, $S_{M}$ instead applies $\mathrm{Eval}$ to $\alpha$ to compute $\{0,1\}^{n}\ni a\mapsto\oplus_{i\in M}a_{i}$ on the encrypted data and returns the new ciphertext $\alpha^{\prime}$. [^3] Here we assume that the QHE scheme allows for more than one homomorphic evaluation, i.e., after given an evaluated ciphertext one can keep evaluating homomorphically any circuit of one’s choice. This property is satisfied by all known QHE schemes. We omit a formal definition of this property and we refer the reader to [9] for details. We also modify step 4 such that if $S$ were to output $b\in\{0,1\}^{n}$, $S_{M}$ instead outputs $\oplus_{i\in M}b_{i}$. Then the following lemma can be proved similarly as [7, Lemma 8] above:
Lemma 5.4.
Theorem 5.5.
[…]
Conversely, there exists a quantum strategy SS and a negligible function $\eta(\lambda)$ such that
[…]
Proof.
By Lemma 5.4, we have
Using this and Eq. 5.1, we obtain that
[…]
@@ 6 Magic square game
by using $(x+y)^{2}\leq 2x^{2}+2y^{2}$, Lemmas 6.6 and 6.5. Similarly, we obtain
@@ Acknowledgments
GM, SS and MW acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2092 CASA - 390781972. GM also acknowledges support by the European Research Council through an ERC Starting Grant (Grant agreement No. 101077455, ObfusQation). MW also acknowledges support by the European Research Council through an ERC Starting Grant (Grant agreement No. 101040907, SYMOPTIC), by the NWO through grant OCENW.KLEIN.267, by the German Federal Ministry of Research, Technology and Space (QuBRA, 13N16135; QuSol, 13N17173), and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, 556164098). MW acknowledges the Simons Institute for the Theory of Computing at UC Berkeley for its hospitality and support while this paper was completed. AM and CP are supported by NSERC Alliance Consortia Quantum grants, reference number number: ALLRP 578455 - 22. AN and CP acknowledge SLMATH for hospitality and financial support during the MIP* = RE hot topics workshop. AN and TZ acknowledge the Simons Institute at UC Berkeley, on whose whiteboard a crucial calculation was performed.
Appendix A Alternate proof of the upper bound
Here we give an alternative proof of the upper bound on the quantum value of compiled XOR game (Theorem 3.1). This alternate approach starts with the following lemma.
Lemma A.1.
$\sqrt{\sum_{s,y\in I_{B}}G_{xs}G_{xy}V_{sy}}=\sum_{y\in I_{B}}G_{xy}\langle u_{x}|v_{y}\rangle.$ (A.1)
Proof.
Then,
since $\|\sum_{s\in I_{B}}G_{xs}v_{s}\|^{2}=\sum_{s,y\in I_{B}}G_{xs}G_{xy}V_{sy}$. Thus we have established Eq. A.1. ∎
The next lemma is proven in the same way as [19, Lemma 26].
Lemma A.2.
Proof.
$(\mathbb{E}(X))^{2}$ $\leq\mathbb{E}(X^{2})$ $\text{for real-valued random variables }X,$ (A.2)
$\left(\sum_{k}p_{k}x_{k}\right)^{2}$ $\leq\sum_{k}p_{k}x_{k}^{2}$ $\text{for }x_{k}\in\mathbb{R},p_{k}\geq 0\text{ and }\sum_{k}p_{k}=1,$ (A.3)
$(\langle\psi|O|\psi\rangle)^{2}$ $\leq\langle\psi|O^{2}|\psi\rangle$ $\text{for quantum states }|\psi\rangle\text{ and observables }O.$ (A.4)
Define $\mathrm{B}_{x}=\sum_{y}G_{xy}B_{y}$. From Eq. A.2, we deduce
Note that the post-measurement states $|\psi_{c\alpha}\rangle$ are not normalized and it holds $\sum_{\alpha}\||\psi_{c\alpha}\rangle\|^{2}=1$. Therefore, using Eq. A.3, we get
from Eq. A.4. ∎
We are now ready to give the alternative proof of our main result.
Alternate proof of Theorem 3.1.
By Lemma A.2, we get
$\beta_{q}(S,\mathcal{G}_{\mathrm{comp}})$ $\leq\sum_{x}\sqrt{\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x)}\sum_{\alpha}\langle\psi_{c\alpha}|\left(\sum_{y}G_{xy}B_{y}\right)^{2}|\psi_{c\alpha}\rangle}.$ (A.5)
[…]
and hence we can upper bound Eq. A.5 as follows:
$=\sum_{x}\sqrt{\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x_{0})}\sum_{\alpha}\langle\psi_{c\alpha}|\mathcal{B}_{x}|\psi_{c\alpha}\rangle}+\eta^{\prime}(\lambda),$ (A.6)
$\sqrt{\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x_{0})}\sum_{\alpha}\langle\psi_{c\alpha}|\mathcal{B}_{x}|\psi_{c\alpha}\rangle}=\sqrt{\sum_{s,y}G_{xs}G_{xy}V_{sy}}=\sum_{y}G_{xy}\langle u_{x}|v_{y}\rangle.$ (A.7)
This follows from Lemma A.1, because $V\in\mathbb{R}^{I_{B}\times I_{B}}$ is positive semidefinite (it is a nonnegative linear combination of Gram matrices) and also satisfies $V_{yy}=1$. As a consequence of this and Tsirelson’s theorem (Theorem 2.10), we obtain that
$\sum_{x}\sqrt{\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x_{0})}\sum_{\alpha}\langle\psi_{c\alpha}|\mathcal{B}_{x}|\psi_{c\alpha}\rangle}=\sum_{x,y}G_{xy}\langle u_{x}|v_{y}\rangle\leq\beta_{q}^{*}(\mathcal{G}).$ (A.8)
Thus we have proved that